# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

### v5

@@ V Experimental Results
In this section, we will report the results obtained in the underwater image enhancement, transmission estimation and color restoration experiments and the ablation study.
V-A Underwater Image Enhancement
We evaluated the proposed PATS-UIENet along with sixteen baselines using three full-reference quantitative metrics and two non-reference quantitative metrics. A qualitative analysis and a performance analysis were also performed. In this subsection, the results obtained in the four experiments will be reported.
V-A1 Full-Reference Quantitative Evaluation
Since the Test-U80, Test-S110 and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR, SSIM [57] and LPIPS [58] metrics across these testing sets. The results are shown in Table I. As can be seen, our PATS-UIENet normally achieved the best performance across all three metrics, regardless of the testing set. Compared with the baselines designed on top of the IFM [1], such as UDCP [4] and DHL [8], and the existing semi-supervised learning approaches, including DDFormer [29], Semi-UIR [28] and UWFormer [30], our method normally showed a large margin.
@@ V-A1 Full-Reference Quantitative Evaluation
TABLE I: The full-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on three real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes
@@
UDCP [4] 9.51 33.66 41.74 10.07 34.29 37.66 14.76 56.43 30.31
DHL [10] 15.16 63.93 30.40 14.61 62.93 28.39 14.99 69.73 22.36
HE [9] 16.60 77.86 25.67 15.41 74.79 30.05 13.61 62.82 36.10
HLRP [15] 13.56 22.76 33.90 12.55 29.24 33.18 12.12 18.21 39.35
MMLE [14] 18.56 76.21 22.57 17.32 77.39 22.77 14.89 62.23 31.99
WaterNet [22] 17.13 70.41 40.01 18.83 75.22 31.08 25.84 84.02 16.28
UColor [23] 21.05 85.24 17.97 20.26 84.25 15.07 24.31 81.40 19.05
FUnIE-GAN [20] 15.31 59.68 44.28 16.89 66.32 34.98 25.11 83.16 13.76
USUIR [24] 16.92 68.69 28.63 17.54 76.58 21.38 15.14 66.51 32.44
PUIE-Net (MP) [25] 22.26 88.89 12.01 21.85 89.19 9.70 21.85 73.47 24.45
U-Transformer [21] 20.98 72.76 26.32 19.92 68.25 30.17 24.09 79.98 15.08
URanker [26] 21.87 85.75 18.64 21.27 87.43 13.51 22.46 82.35 20.87
CCMSRNet [27] 22.74 88.74 13.18 22.22 89.07 10.76 18.96 77.67 23.85
DDFormer [29] 10.94 25.74 58.26 10.87 41.27 72.19 10.64 27.81 69.46
Semi-UIR [28] 23.63 81.81 23.08 19.83 73.95 30.54 19.12 74.94 25.68
UWFormer [30] 19.65 85.26 17.21 21.45 90.27 10.88 19.62 80.36 21.05
PATS-UIENet (Ours) 23.59 90.16 10.42 22.78 90.54 8.85 25.87 84.34 12.04
V-A2 Non-reference Quantitative Evaluation
We used two non-reference metrics, i.e., UIF [59] and MUSIQ [60], to assess the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets. The results are shown in Table II. It can be seen that our method achieved the best UIF score on three out of the six testing sets, including Test-U80, Test-S110 and Test-C60, and ranked the second on the Test-R53 testing set. Regarding the MUSIQ metric, the proposed method still demonstrated the competitive performance, especially on the challenging Test-S110, Test-Scenes and Test-R53 testing sets, even though it did not always produce the highest score.
TABLE II: The non-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on six real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes Test-C60 Test-UCCS Test-R53
[…]
UDCP [4] 36.40 44.99 32.01 51.38 61.25 36.99 25.41 33.03 34.72 29.72 3.24 36.85
DHL [10] 41.32 47.43 14.80 57.22 0.55 36.57 25.04 37.27 3.62 31.07 2.53 42.64
HE [9] 44.04 46.68 45.01 56.14 52.81 35.79 41.50 36.08 38.08 31.24 1.67 43.06
HLRP [15] 0.59 49.73 0.91 56.87 0.71 40.21 3.57 34.65 0.32 34.44 1.26 43.53
MMLE [14] 30.99 52.50 35.05 60.18 42.05 44.67 24.95 40.18 28.08 35.69 24.67 53.53
WaterNet [22] 33.25 31.84 28.50 40.47 72.94 38.47 28.65 34.75 33.38 24.88 9.71 32.93
UColor [23] 44.15 45.10 42.83 54.80 70.20 35.12 26.46 36.01 49.22 35.12 2.31 49.63
FUnIE-GAN [20] 29.12 47.47 22.77 52.27 37.69 45.38 23.11 43.49 22.67 34.64 10.93 38.92
USUIR [24] 5.17 44.44 37.02 56.83 48.82 37.91 22.35 34.15 20.54 28.83 17.98 39.44
PUIE-Net (MP) [25] 9.18 49.05 60.98 60.44 27.69 45.46 3.39 38.69 51.76 29.18 46.16 45.15
U-Transformer [21] 5.16 39.34 24.94 44.96 38.68 30.44 29.31 37.66 29.69 27.05 19.24 30.06
URanker [26] 54.87 44.63 58.31 57.15 72.54 39.20 44.70 38.37 45.97 29.60 39.19 41.92
CCMSRNet [27] 54.61 49.85 0.23 60.59 63.59 39.64 41.85 40.24 51.34 33.72 34.80 48.23
DDFormer [29] -5.36 36.99 1.92 40.68 -3.16 23.87 11.14 36.21 -2.14 20.19 5.86 24.87
Semi-UIR [28] 50.06 45.40 31.94 51.22 44.01 41.47 46.82 43.25 32.28 30.34 23.33 35.21
UWFormer [30] 46.48 50.95 55.69 61.53 67.40 36.80 36.76 41.40 53.24 34.30 39.98 46.26
PATS-UIENet (Ours) 58.83 49.91 61.20 60.66 68.07 46.70 46.99 40.49 51.18 31.63 42.51 46.63
V-A3 Qualitative Analysis
The enhanced images generated by the 16 baselines and our method on six testing sets are shown in Figs. 4 - 9, respectively. It can be seen that some methods with the higher UIF [59] or MUSIQ [60] score did not produce visually pleasing results. For example, MMLE [14] produced the highest MUSIQ score on Test-U80 and Test-R53, but the resulting images suffered from pale colors and over-enhancement artifacts (see Figs. 4 and 9). In contrast, our PATS-UIENet improved different degraded images and produced images with the natural and vivid color. Although the UIF [59] or MUSIQ [60] scores that our method produced were slightly lower than those obtained using some baseline methods [14, 22, 20] on certain testing sets, the enhanced images still manifest visually satisfactory quality.
Figure 4: The results produced by 16 baselines and our method in terms of a degraded image in the Test-U80 testing set.
Figure 5: The results produced by 16 baselines and our method in terms of a degraded image in the Test-S110 testing set.
Figure 6: The results produced by 16 baselines and our method in terms of a degraded image in the Test-Scenes testing set.
Figure 7: The results produced by 16 baselines and our method in terms of a degraded image in the Test-C60 testing set.
Figure 8: The results produced by 16 baselines and our method in terms of a degraded image in the Test-UCCS testing set.
Figure 9: The results produced by 16 baselines and our method in terms of a degraded image in the Test-R53 testing set.
V-A4 Performance Analysis
We compared the proposed PATS-UIENet with 11 deep baseline methods in terms of the number of parameters, Floating Point Operations Per Second (FLOPs) and inference speed (ms). Both the number of parameters and the FLOPs were calculated using the ptflops [^2] https://pypi.org/project/ptflops/ tool. As shown in Table III, the values of number of parameters, FLOPs and inference speed vary significantly. Specifically, UColor [23] exhibits the largest number of parameters (148.04M) and the largest FLOPs value (2804.36G), while UWFormer [30] incurs the slowest inference speed (491.00 ms). On the other hand, USUIR [24] has the lightest design with only 0.23M parameters and has a FLOPs value of 29.62G and owns the fastest inference speed (2.13ms). In contrast, our PATS-UIENet achieved a proper balance between the number of parameters or FLOPs and the inference speed.
TABLE III: Comparison of between 11 deep UIE methods and our approach in terms of number of parameters, FLOPs and inference speed.
Method #Params (M) FLOPs (G) Speed (ms)
WaterNet [22] 1.09 142.84 6.40
UColor [23] 148.04 2804.36 15.51
FUnIE-GAN [20] 7.02 20.48 73.18
USUIR [24] 0.23 29.62 2.13
PUIE-Net (MP) [25] 1.40 70.54 101.89
U-Transformer [21] 22.8 5.96 91.30
URanker [26] 3.15 20.90 45.50
CCMSRNet [27] 21.69 87.18 152.51
DDFormer [29] 7.63 35.54 50.04
Semi-UIR [28] 1.65 72.88 84.93
UWFormer [30] 29.84 30.18 491.00
PATS-UIENet (Ours) 45.71 179.70 86.17
V-B Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table IV. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [39]. As shown in Fig. 10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
TABLE IV: The quantitative evaluation of four prior-based baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration, respectively.
@@ V-B Transmission Estimation and Color Restoration
DCP [3] -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
UDCP [4] -0.54 36.97 -0.15 40.43 0.06 38.80 0.12 51.45
GDCP [5] 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
DHL [10] -0.07 32.10 0.10 35.50 0.36 35.92 0.11 34.82
Ours 0.62 12.83 0.32 9.47 -0.28 15.31 0.04 13.14
V-C Ablation Study
To investigate the effect of different components of the PATS-UIENet, we conducted a series of ablation experiments. For simplicity, we only utilized Test-U80 and the Michmorest subset in Test-R53.
V-C1 Effect of the Semi-supervised Learning Framework
To validate the effect of the proposed semi-supervised learning framework, we compared it with the unsupervised learning, supervised learning and bi-directional supervised learning frameworks. As reported in Table V, the proposed semi-supervised framework always outperformed the other frameworks in terms of different metrics across different testing sets. It has been suggested that our network can be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improves the enhancement performance.
TABLE V: Comparison of different learning frameworks on Test-U80 and the Michmorest subset in Test-R53.
Learning Framework Test-U80 Michmoret
@@ V-C1 Effect of the Semi-supervised Learning Framework
Unsupervised 16.85 78.44 29.91 0.50 5.46
Supervised 22.19 88.23 13.07 0.57 17.01
Bi-supervised 23.00 89.49 11.93 0.61 15.02
Semi-supervised (Ours) 23.59 90.16 10.42 0.62 12.83
V-C2 Effect of the RCT and RCM
To examine the effect of the RCT and RCM on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them. In this case, we derived three variants of the PATS-UIENet: one with both the RCT and RCM removed, one with only the RCT removed and one with only the RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively. As shown in Table VI, the complete PATS-UIENet achieved the best performance in terms of all the metrics. Removal of the RCT led to a slight increase in the PSNR and SSIM metrics compared to the Simplified version, but worsened the LPIPS and $\bar{\psi}$ scores. On the other hand, removal of the RCM reduced the performance with regard to all metrics across the two data sets, confirming its usefulness in facilitating information exchange between the B-Stream and A-Stream. The results demonstrate that both the RCT and RCM contribute meaningfully to the performance of our method.
TABLE VI: Impact of the RCT and RCM on the performance of our method on Test-U80 and the Michmorest subset in Test-R53.
Variant Test-U80 Michmoret
[…]
Simplified 22.96 89.64 11.25 0.59 13.59
w/o RCM 22.10 88.83 12.20 0.60 13.50
w/o RCT 23.14 90.03 12.20 0.60 13.97
Ours 23.59 90.16 10.42 0.62 12.83
V-C3 Effect of the Degradation Control Factor
To evaluate the impact of the degradation control factor $\alpha$ on the performance of the PATS-UIENet, we conducted an ablation experiment on testing three different values of $\alpha$. As shown in Table VII, the default setting of $\alpha$ = 0.1 achieved the best performance with regard to both the full-reference and the non-reference metrics across the two data sets.
TABLE VII: The effect of the degradation control factor, $\alpha$, on the performance of our method.
$\alpha$ Test-U80 Michmoret
[…]
0.001 23.21 89.77 11.02 0.59 13.92
1 23.06 89.72 11.08 0.61 14.11
0.1 (Ours) 23.59 90.16 10.42 0.62 12.83
V-C4 Effect of Loss Hyperparameters
To evaluate the impact of the hyperparameters used for the loss functions, including $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{unsup}$, on the performance of the proposed PATS-UIENet, we conducted a comprehensive ablation study by changing their values. Regarding the supervised learning loss function (see Eq. (7)), the results produced by our method with different combinations of the $\lambda_{1}$ and $\lambda_{2}$ values are shown in Table VIII. It can be seen that the combination that we used produced the best result in terms of each metric. For the unsupervised learning loss function (see Eq. (15)), the results produced by our method with different values of $\lambda_{3}$ are reported in Table IX. Again, the value that we chose led to the best result no matter what metric was considered. In terms of the semi-supervised learning loss function (see Eq. (16)), we present the results obtained using our method with different $\lambda_{unsup}$ values in Table X. As can be observed, the value that we utilized produced the best result with regard to each metric. The above findings highlight the effectiveness of the values of different loss hyperparameters that we chose.
TABLE VIII: The effect of different combinations of hyperparameters used for the loss function of the supervised learning scheme, i.e., $\lambda_{1}$ and $\lambda_{2}$, on the performance of our method.
$\lambda_{1}$ $\lambda_{2}$ Test-U80 Michmoret
[…]
1 0.0005 22.04 88.01 12.94 0.36 16.92
0.0001 0.0005 22.84 89.65 10.85 0.61 13.57
0.1 0.1 22.95 89.67 11.52 0.61 15.69
0.1 0 22.91 80.57 11.54 0.59 16.92
0.1 0.0005 (Ours) 23.59 90.16 10.42 0.62 12.83
TABLE IX: The effect of the weighting factor used for the loss function of the unsupervised learning scheme, i.e., $\lambda_{3}$, on the performance of our method.
$\lambda_{3}$ Test-U80 Michmoret
[…]
0.1 22.95 89.80 11.01 0.61 15.14
10 22.36 88.19 13.09 0.60 14.72
1 (Ours) 23.59 90.16 10.42 0.62 12.83
TABLE X: The effect of the weighting factor used for the loss function of the semi-supervised learning scheme, i.e., $\lambda_{unsup}$, on the performance of our method.
$\lambda_{unsup}$ Test-U80 Michmoret
[…]
1 22.32 88.75 12.16 0.61 14.08
0.01 23.31 89.96 10.92 0.61 12.84
0.1 (Ours) 23.59 90.16 10.42 0.62 12.83
V-C5 Effect of Different Stream Architectures
To further examine the effect of different stream architectures on the PATS-UIENet, we constructed its three variants by building the D-Stream, B-Stream and A-Stream using convolution and/or Transformer networks. In addition, we obtained a fourth variant by removing the D-Stream. In essence, this variant is equal to the network built on top of the original IFM [1]. The four variants were trained and tested using the same setup as that used for our PATS-UIENet. The results produced by these variants and our method are shown in Table XI. It can be seen that our method produced the better, or at least comparable, results in contrast to the four variants. This finding supports the design philosophy of the PATS-UIENet. That is to say, CNNs are suitable for capturing local characteristics while Transformer is good at encoding global characteristics. Moreover, our network which was built on top of the revised IFM [2] normally outperformed the variant constructed based on the original IFM [1], confirming the choice of the revised IFM [2].
TABLE XI: Comparison between our method and its four variants obtained by building the D-Stream, B-Stream and A-Stream using different networks.
Method Streams Test-U80 Michmoret
[…]
$\text{Variant}_{1}$ CNN CNN CNN 22.31 88.69 12.48 0.58 12.44
$\text{Variant}_{2}$ Trans Trans CNN 20.71 75.36 25.31 0.73 17.28
$\text{Variant}_{3}$ Trans Trans Trans 20.57 75.00 25.73 0.71 14.23
$\text{Variant}_{4}$ N/A CNN Trans 22.51 88.85 11.47 0.71 15.95
Ours CNN CNN Trans 23.59 90.16 10.42 0.62 12.83
VI Conclusion
@@
In this paper, we introduced a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, namely, PATS-UIENet, which explicitly estimates the degradation parameters of the revised IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first effort to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, sixteen baselines on six underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.

### v6

@@ 5 Experimental Results
In this section, we will report the results obtained in the underwater image enhancement evaluation, parameter estimation and color restoration evaluation and ablation study.
5.1 Evaluation on Underwater Image Enhancement
The quantitative evaluation of eight baselines and the proposed method was conducted for the UIE task. The results are reported in Tables 1. As can be seen, our method outperformed both six prior-based methods and two supervised deep learning approaches on two labeled data sets, including Test-U80 and Test-S110, when the PSNR and SSIM metrics were used. On the other hand, the UIQM and UCIQE metrics were also used in the evaluation experiment where we did not utilize reference images. In this case, the three deep learning methods were inferior to the prior-based approaches. This finding is consistent with those reported in the existing studies [16, 18]. Specifically, the performance of our method was normally comparable to that of the two deep learning approaches, i.e., WaterNet [16] and UColor [17], and was comparable to that of the prior-based methods. As demonstrated in [16, 18, 25], both the UIQM and UCIQE metrics are not always consistent with the quality of the images processed by UIE methods. It should be noted that, however, the prior-based methods did not produce stable results across the five testing sets. This should be attributed to the fact that these methods normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene is not suited to the prior assumption [4, 16].
@@ 5.1 Evaluation on Underwater Image Enhancement
Table 1: The quantitative evaluation of eight baselines and our method on five real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts respectively.
Method Test-U80 Test-S110 Test-C60 Test-UCCS Test-R53
@@
DCP 12.48 63.34 1.35 0.51 14.52 67.38 1.36 0.57 0.56 0.51 0.03 0.44 -2.05 0.42
UDCP 9.37 33.62 2.69 0.59 10.04 34.28 1.23 0.60 1.24 0.54 1.83 0.57 0.76 0.59
GDCP 12.46 67.41 2.17 0.56 14.67 71.29 2.12 0.64 1.49 0.60 1.24 0.56 5.50 0.50
DHL 14.49 62.81 3.19 0.64 14.96 69.68 2.35 0.66 1.85 0.64 2.98 0.63 0.22 0.64
HE 16.48 77.77 3.58 0.67 15.37 74.76 2.97 0.68 2.32 0.65 3.93 0.67 1.71 0.64
Retinex 1038 61.25 2.33 0.48 8.29 53.44 0.85 0.45 1.21 0.46 2.03 0.47 -2.03 0.51
WaterNet 17.02 70.27 2.11 0.50 18.80 75.19 1.47 0.54 1.28 0.51 1.25 0.45 0.09 0.45
UColor 20.93 85.19 3.44 0.60 20.23 84.23 2.30 0.63 2.08 0.58 3.22 0.55 1.28 0.54
Ours 22.82 88.87 3.60 0.57 20.46 87.22 2.35 0.60 2.02 0.56 3.24 0.53 1.92 0.56
The qualitative comparison of the results produced by the eight baselines and our method is shown in Fig.4. It can be seen that the enhanced images derived using our method presented the better visual quality, compared with the images enhanced by the baselines. Our method not only effectively removed the color casts in the degraded images but also greatly increased the visibility of them. In contrast, the prior-based methods usually introduced artifacts that decreased the visual quality while the learning-based methods could not eliminate the color casts completely.
Figure 4: The results obtained using eight baselines and our method on five degraded images in Test-U80, Test-S110, Test-C60, Test-UCCS and Test-R53, respectively, shown from top to bottom.
5.2 Evaluation on Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and eight baselines for transmission estimation and color restoration is reported in Table 5.2. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the eight baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets no matter which metric was considered, while it performed comparably to the best baseline on the Nachsholim and Satil subsets. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [25]. As shown in Fig. 5, our method is able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
Table 2: The quantitative evaluation of eight baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration respectively.
@@ 5.2 Evaluation on Transmission Estimation and Color Restoration
DCP -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
UDCP -0.54 36.97 -0.15 40.43 0.06 38.79 0.12 51.45
GDCP 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
DHL -0.07 32.10 0.10 35.50 0.39 35.91 0.11 34.82
HE - 35.44 - 35.43 - 34.37 - 36.33
Retinex - 31.44 - 34.37 - 33.29 - 35.31
WaterNet - 25.40 - 25.70 - 25.80 - 28.31
UColor - 15.55 - 13.94 - 17.21 - 21.43
Ours 0.71 11.51 0.48 8.75 0.18 19.59 0.04 13.46
Table 3: The ablation study on Test-U80 and the Michmorest subset in Test-R53.
Method Test-U80 Michmoret
@@
Simplified 22.16 89.57 -0.15
w/ RCT 22.97 89.93 0.58
w/ Bi-Sup 22.51 88.94 0.68
w/ Semi-Sup 22.82 88.87 0.71
5.3 Ablation Study
To examine the effect of each component of our method, including the RCT, bi-directional supervised scheme and semi-supervised learning framework, we conducted an ablation study. For simplicity, we only utilized Test-U80 and the Michmoret subset in Test-R53. The PA-UIENet without the RCT together with the forward-enhancement supervised learning module was crafted as a simplified version. On top this version, we derived three additional variants by inserting the RCT (w/ RCT) and then replacing the forward-enhancement supervised module by the bi-directional supervised scheme (w/ Bi-Sup) or the semi-supervised learning framework (w/ Semi-Sup). As shown in Table 5.2, the simplified version outperformed all the baselines in terms of the PSNR and SSIM metrics. However, this variant did not estimate transmission maps well, which may be attributed to the overfitting problem resulted from limited training data. When the RCT was used, the performance of image enhancement and transmission estimation was boosted. It was indicated that the adaptability of the red channel increased the robustness and generalization of the model. Though the utilization of the bi-directional-supervised learning scheme or the semi-supervised learning framework had a negative effect on the PSNR and SSIM values, the effectiveness of transmission estimation continuously increased. It is suggested that the use of the backward-degradation learning module or the unsupervised learning scheme alleviates the overfitting problem.
6 Conclusion
@@
In this paper, we developed a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet, which explicitly estimates the degradation parameters of the IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first attempt to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, eight baselines on five real-world underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.